Sam is looking to buy his first flat, and has R15 000 in cash savings which he will use as a deposit. He has viewed a flat which is on the market for R250 000, and he would like to work out how much the monthly repayments would be. He will be taking out a 30 year mortgage with monthly repayments. The annual interest rate is 11%.


1
Expert's answer
2021-07-08T12:06:23-0400

Principal mortgage amount (P) : 250,000 - 15,000 = 235,000


Annual interest rate (i) = 11%


Loan term = 30 years


Number of monthly payments (n) = "30 \times 12= 360"


Monthly mortgage payment (M) = "\frac{P [ i (1 + i) ^n]}{[(1+i) ^n -1]}"


i= "\frac{11}{12}" = 0.9167


M= "235000[ 0.9167( 1+ 0.9167)^{360}]\/ [(1+0.9167)^{360} -1]"


M= 2,237.96


Therefore monthly repayments = R2,237.96